OpenStudy (anonymous):

A satellite completes one orbit in one day (86,400 seconds). The orbital distance is 22,500 kilometers. What is the average velocity of the satellite at the end of one orbit?

OpenStudy (anonymous):

we have formula to find out the velocity: \(v= \dfrac{2\pi r}{T}\) where T is period =86400seconds and r = 2.25 *10^7